OpenStudy (anonymous):

Solve. 0.3r + 0.2 = 1.7

OpenStudy (anonymous):

\[\frac{3}{10}r + \frac{2}{10} = \frac{17}{10}\] multiplying the tops and bottoms by 10 makes this alot easier to do

OpenStudy (anonymous):

sorry, multiplying the top, dividing the bottom